OpenStudy (anonymous):

3y+x=3 -2y+5x=15 Help? (3,0)?

OpenStudy (anonymous):

yes.

OpenStudy (gw2011):

To verify if your answer is correct here are the steps in solving this: You want to eliminate one of the unknowns - in this case I chose x. So, multiply one of the equations by -5 to get: -15y-5x=-15 (this first equation was multiplied by -5) -2y+5x=15 Adding these two equations you get: -17y=0 y=(0)/(-17) y=0 (now substitute this value into one of the equations - I chose the first one) 3y+x=3 (3)(0)+x=3 x=3 So, the solution is x=3 and y=0 which is (3,0)